OpenStudy (anonymous):

solve the equation. check your answer 3x-7=5x+19?

mathslover (mathslover):

\[\large{3x-7=5x+19}\] \[\large{3x-3x-7-19=5x+19-3x-19}\] \[\large{-26=2x}\] \[\large{x=\frac{-26}{2}}\] \[\large{x=-13}\]

mathslover (mathslover):

This is better ?

mathslover (mathslover):

did u get it @pacemlpace

OpenStudy (anonymous):

friend add 7 to both sides and subtract 5x to both sides then you will get your answer eeasily.